OpenStudy (anonymous):

Which Scenario describes a situation where x varies directly with y? a. The difference between x and y is a constant. b. The sum of x and y is a constant. c. The product of x and y is constant. d. The quotient of x and y is constant.

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

:( help pleaseee

OpenStudy (igreen):

I believe it refers to multiplication.

OpenStudy (igreen):

So what is your answer going to be? @Mimi_Styles

OpenStudy (mathmath333):

the situation can be represented by this equation \(y ~\alpha ~x\\~\\ y=kx\)

OpenStudy (igreen):

'x varies directly with y' Direct Variation is stated \(y = kx\) Where 'k' is the constant.

OpenStudy (anonymous):

\(x~\alpha ~~y\\x=ay \\ \dfrac{x}{y}=a\)

OpenStudy (mathmath333):

ie \( \huge \dfrac{y}{x}=constant\)

OpenStudy (anonymous):

i gave u a hint xD

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

i saw all comments at time lol

OpenStudy (anonymous):

yeah that's why I got confused lol. let me read them all

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

okay I read it. but still don't really know what the answer would? like would it be ? like would it be c?

OpenStudy (anonymous):

@Marki come backkkk D: @iGreen

OpenStudy (anonymous):

naa try again

OpenStudy (anonymous):

d? D:

OpenStudy (anonymous):

I hate math ._. :(

OpenStudy (anonymous):

i love math and yes ur right ;)

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

I fanned both of ya'll :)

OpenStudy (anonymous):

can I ask three more questions? :3

OpenStudy (anonymous):

:3 k

OpenStudy (anonymous):

yay xD

OpenStudy (igreen):

Yep..D is correct..

OpenStudy (anonymous):

10. A line with a slope of -3 passes through (2,5) and (4,d). What is the value of d? a. -1 b. 2 c. 8 d. 11

OpenStudy (igreen):

Plug in what you know into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(-3 = \dfrac{d-5}{4-2}\) Plug in -1, 2, 8, and 11 for 'd' and see which one comes up with '-3'.

OpenStudy (anonymous):

how do I plug it in? I don't have a calculator :(

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

is there a site I can use a calculator on? :((

OpenStudy (anonymous):

@Marki @iGreen guysssssssssssssssssssssssssssssssssssssss :((

OpenStudy (anonymous):

anyoneeee out thereeee

OpenStudy (anonymous):

yessss but what kind of calculator ? like I need the right one :(

OpenStudy (anonymous):

oh lawd

OpenStudy (igreen):

Lol

OpenStudy (anonymous):

lmao.

OpenStudy (anonymous):

wait there is much cool sites for calculator

OpenStudy (anonymous):

http://web2.0calc.com/ I've been using this but I don't know how to put it in :/

OpenStudy (anonymous):

I ain't smart. -knocks on my head- see? nothing

OpenStudy (anonymous):

https://www.desmos.com/calculator/p71mx55qkg

OpenStudy (anonymous):

I'm not*

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

both are good save them and use them :)

OpenStudy (anonymous):

yass Desmossssss I can use desmos to find the answer?

OpenStudy (anonymous):

okay. wait

OpenStudy (igreen):

^^^^^^^^^^^^^^

OpenStudy (anonymous):

wolfram helps u more demoss is like hand graphing calculator

OpenStudy (igreen):

O: CONGRATZ ON 99! :O @Nnesha

OpenStudy (anonymous):

>_>

OpenStudy (igreen):

Just plug in a number..it's just simple subtraction & division.

OpenStudy (igreen):

What's 4 - 2?

OpenStudy (anonymous):

2 -_-

OpenStudy (anonymous):

-huffs and does it-

OpenStudy (igreen):

Yes, so we can simplify our equation to: \(-3 = \dfrac{d - 5}{2}\) Now it's much easier.

OpenStudy (anonymous):

oh? okay hold on

OpenStudy (anonymous):

uhhhh

OpenStudy (anonymous):

well peps, I guess I'm going to be working at a hotdog stand :(

OpenStudy (anonymous):

oh wait nvm I think I did this all wrong hold on

OpenStudy (anonymous):

it's A

OpenStudy (anonymous):

right??

OpenStudy (anonymous):

yayyyyyyyyyyyyyy! *throws my shades on* B-)

OpenStudy (igreen):

Yep, you got it! :D